# KiDS-Legacy: angular galaxy clustering from deep surveys with complex selection effects

### v1

@@ 2.2 The measurement of two-point correlation function
In practice, the galaxy 2PCF is measured in angular bins by counting the number of galaxy pairs (one galaxy from catalogue $a$ and the other from $b$) with separation angles within each angular bin. Given a pair of galaxy catalogues (denoted as $D$ hereafter) and corresponding random catalogues (denoted as $R$ hereafter), the 2PCF can be measured using a standard Landy & Szalay (1993) estimator:
@@
where $\theta_{i}$ is the mean angle in the $i$-th angular bin. $(D^{a}D^{b})_{i}$ is the normalized number of galaxy pairs with separation angles within the $i$-th bin, which can be formally expressed as
@@
where $N_{D}$ is the number of galaxies, $\theta_{pq}$ is the separation angle between the $p$-th and $q-$th galaxies, $\Theta_{i}(\theta)$ is the rect function that equals 1 when $\theta$ falls in the $i-$th bin and 0 otherwise. $\delta_{\mathrm{K}}^{ab}$ is the Kronecker $\delta$ symbol. The denominator in Eq. (8) is the total number of galaxy pairs and the $\delta^{ab}$ term concerns removing galaxies paired with themselves for auto-correlation (i.e., $a=b$). The $RR$ term is likewise defined while $DR$ is defined without $\delta_{\mathrm{K}}^{ab}$.
In this work, instead of counting galaxy pairs in angular bins, we pixelise the galaxy distribution into the Healpix scheme (Gorski et al., 2005) and count pixel pairs weighted by galaxy numbers in each pixel. This will lose sub-pixel information, but we have three reasons for doing this: 1) pixelisation can speed up the calculation: the original KiDS catalogue contains $\sim 200$ million galaxies, while a pixelised galaxy map with Nside=2048 which we will use throughout this work only has $\sim 1$ million non-zero pixels, so counting pixel pairs instead of galaxy pairs will significantly speed up the measurement; 2) as will be introduced later, our method to recover ORs is pixel-based, so any sub-pixel variable depth cannot be corrected; 3) the pixel size for an Nside=2048 HEALPix map is 1.7 arcmin, corresponding to a physical size of 0.86 Mpc at $z=1.5$, which is much below the linear scales that we will probe in this work.
@@
The random catalogue or map is used to factor out galaxy distributions induced by cosmology-unrelated systematics, including galactic and atmospheric foregrounds, instrument response, and survey strategy. If all those systematics are uniform across the survey footprint, the random catalogue is usually taken as a uniform Poisson random sample within the survey footprint with several times the mean galaxy number density. However, deep surveys, such as the KiDS survey, have non-uniform systematics that cause variable depth in the galaxy sample. A UR will fail to capture the spatial variation of galaxy number density caused by those selection effects and biases the 2PCF.
To de-bias the 2PCF, we can create a random catalogue with the same variable depth, namely an ‘organised random’ catalogue, or, for pixelised measurement, an OR weight map for the $R$ terms in the estimator. The method to recover the OR will be presented in Section. 4. We note that the OR can be used for summary statistics other than the angular 2PCF. We will present the OR application in angular power spectra $C_{\ell}$ in Appendix B.
@@ 3.1 The Kilo-degree Survey and the Legacy data
Figure 2: Galaxy distribution of the KiDS-Legacy catalogue. The map is pixelised into HEALPix grids with $\texttt{Nside}=2048$. The map is colour-coded by galaxy number per pixel, which has a size of 1.7 arcmin. A tile-based selection pattern can be seen by eyes.
Table 1: A summary of systematics that we use to recover the OR. The second column indicates the spatial variance type: inter-tile (type-A), intra-tile (type-B), and tile-independent (type-C). The other columns indicate their definitions, physical origins, and how they potentially select the galaxy sample.
@@
In this work, we will use the KiDS-Legacy galaxy catalogue selected from the fifth data release (DR5) of the Kilo-degree Survey (KiDS, Wright et al., 2024). KiDS is a wide-field imaging survey that measures the positions and shapes of galaxies using the VLT Survey Telescope (VST) at the European Southern Observatory (ESO). Both the telescope and the survey were primarily designed for weak gravitational lensing applications. High-quality optical images are produced with VST-OmegaCAM, and these data are then combined with imaging from the VISTA Kilo-degree INfrared Galaxy survey (VIKING; Edge et al. 2013), allowing all sources in KiDS to be photometrically measured in nine optical and near-infrared bands: $ugriZYJHK_{\mathrm{s}}$ (Wright et al., 2019). Although the sky coverage of KiDS is smaller than some galaxy lensing surveys (such as DES, Abbott et al., 2016), galaxy photometric redshift estimation and redshift distribution calibration (especially at high redshift) benefit from complementary NIR information from VIKING (which was co-designed with KiDS to reach complementary depths in the NIR bands). Cosmological constraints have already been made available from DR3 and DR4 (Hildebrandt et al., 2017; Heymans et al., 2021). KiDS DR5 covers a survey area of 1347 $\mathrm{deg}^{2}$. It also includes an $i$-band re-observation of the full footprint, thereby increasing the effective $i$-band depth by 0.4 magnitudes and enabling multi-epoch science. There is a 27 $\mathrm{deg}^{2}$ overlap with deep spectroscopic surveys, which enables the robust calibration of photometric redshifts across the full survey footprint (Hildebrandt et al., 2021).
@@
The “KiDS-Legacy” sample used in this work is a subset of the full DR5 sample primarily determined by the availability of reliable shape measurements. It is the lensing sample that will be used for the fiducial KiDS DR5 cosmological analyses. Each galaxy in the KiDS-Legacy sample has ellipticities measured with the lensfit algorithm (Miller et al., 2013; Fenech Conti et al., 2017; Li et al., 2023a) for weak lensing analyses. The footprint of the KiDS DR5 is divided into a Northern and a Southern patch. Artefacts around bright, saturated starlight, planets, the Moon, satellite flares, aeroplanes, and higher-order reflections from very bright stars have been masked out, leaving an unmasked $\sim 1000\mathrm{deg}^{2}$ footprint. In addition, blended sources, unresolved binaries, transients, point sources flagged by lensfit, and sources with failed photometry, badly estimated shapes, low resolution, or zero lensing weight are removed from the full sample, resulting in a Legacy sample of 43,205,156 galaxies, corresponding to 12 galaxies per arcmin-2. A Bayesian photo-$z$ (BPZ, Benítez, 2000) estimation gives an approximated redshift range of $2.00.1<z_{\mathrm{B}}<$ 2.0. For more details about the construction of the KiDS-Legacy catalogue, we direct the reader to section 7 of Wright et al. (2024). The galaxy distribution is shown in Figure 2. We can tell by eye that the galaxy distribution has a clear tile-like pattern that is not likely the LSS.
@@ 3. Type C,
A synthesis of systematics results in a complex selection pattern, namely variable depth, in the galaxy distribution. Figure 2 illustrates the galaxy map of the KiDS-Legacy catalogue, which exhibits a tile-based pattern of variable depth. For example, galaxies in the anomalous dark blue tile around ($\alpha=196\deg$, $\delta=1.5\deg$) in the centre of KiDS-North in Figure 2 are significantly depleted. Heydenreich et al. (2020) studied the effect of variable depth on cosmic shear and concluded that the impact is insignificant for KiDS-like surveys. However, the variable depth will introduce significant bias in angular clustering measurement if not corrected. As galaxy clustering measurements form part of the KiDS-Legacy $3\times 2$pt and $6\times 2$pt analyses, it is essential to correct this bias.
@@
To correct the selection effects, we create an OR which reflects the same variable depth as the galaxy sample from the systematics given by the KiDS-Legacy catalogue. Those systematics generally originate from the Galactic foreground, atmospheric seeing, survey strategy, and instrument setup, and are ideally independent of cosmology, galaxy properties, and the LSS. In the catalogue, systematics are described by metadata that are calculated or derived from the instrument setup, observing conditions, image properties, and so on. In the following text, we will use “systematics” to also refer to the metadata we use to obtain the OR. We select five systematics that are most likely to have selection effects: {Level, PSF_size, PSF_ell, EXTINCTION_r, GAIA_nstar}. Their types of distribution, definitions, origins and selection effects are summarised in Table 1. These systematics, with the exception of GAIA_nstar, are measured in the $r$-band because this is the detection band and we do not consider tomographic redshift binning in this analysis.
@@
Figure 3 shows the maps of these systematics. Each systematics map is divided into south and north fields, colour-coded by the mean systematics value in each pixel. The black curves in the colour bars are the probability distributions of the systematics. From Figure 3, we can see how Level, PSF_size, PSF_ell are distributed based on the $1^{\circ}\times 1^{\circ}$ tiles. Among them, Level and PSF_size are determined by the seeing of each exposure and are therefore relatively uniform in each tile, but vary from tile to tile; while PSF_ell depends on the curvature of the focal plane, which changes from the centre to the edge of each tile. On the other hand, EXTINCTION_r [^5] Given that the extinction in different bands is commonly just assumed to be the same template scaled by a different factor (thus they are fully correlated), we only use the $r$-band extinction for our model training. and Gaia_nstar reflect the spatial distribution of the Milky Way and are therefore more diffuse and independent of the tiles.
@@
Note that we could recover the OR with all the systematics such as PSF shape components and background counts, but in practice, it is advantageous to only use a subset of systematics that is most likely to select galaxies to reduce computation time. The rest of the systematics are either not correlated with galaxy numbers or are strongly correlated with the selected systematics. In Appendix C, we run our method with all the systematics provided in the catalogue and prove that the performance does not improve. In the following sections of this paper, we present and test the method for recovering the organised random from these five systematics.
@@
Figure 3: The KiDS-Legacy systematic maps considered in this paper. Each systematics map is divided into northern and southern fields, plotted together with their colour bars. The black curves in the colour bars show the probability distributions of each systematics.
@@ 4 The SOM+HC method to recover organised randoms
Figure 4: The flowchart of the measurement of the 2PCF with reconstructed organised randoms. The colour block after the 2PCF indicates the colours shown in the following $w(\theta)-\theta$ figures.
Figure 4 shows a flowchart of the 2PCF measurement of galaxies with variable depth. Through observation, we obtain a depleted catalogue of galaxies with systematics evaluated for each galaxy. If we follow the upper half of the flowchart by using UR to compute the 2PCF, we get a biased “UR” 2PCF. Instead, we can recover OR to correct for non-uniform selection effects and obtain the unbiased “OR” 2PCF.
@@ 4.1 The selection effects induced by systematics
If the systematics are uncorrelated, the overall selection function is a multiplication of the selection functions of individual systematics. One can reconstruct the total selection function by modelling the selection effect of each systematics (Rodríguez-Monroy et al., 2022). In practice, however, it is not always possible to construct reliable quantitative models to describe the effect of each systematics on the galaxy number density. In addition, systematics can be correlated (e.g. extinction in different bands, extinction and GAIA stars, etc.), making the overall selection function more complicated. In this work, we use an unsupervised machine learning method, namely SOM+HC to recover the OR weight to account for the selection effect. This method does not assume formal models for selection functions and their correlations, so it is flexible enough to account for arbitrarily complex selection effects.
@@ 4.2 Self-organising map and hierarchical clustering
In summary, recovering the ORs amounts to identifying the LSS-independent clustering of galaxies according to their systematics and to finding the effective sky area occupied by galaxies from each cluster. Assuming that galaxies from each cluster have the same selection, we generate a UR in each effective sky area and then combine those random samples from all the clusters. Therefore, the key is to group galaxies in systematic space. Hierarchical clustering (Murtagh & Contreras, 2012) is a widely-used clustering algorithm. It has the advantages of flexibility, robustness, and interpretability, but is computationally expensive (the computational time scales with the data volume as $O(N^{2})$, where $N$ is the number of galaxies). On the other hand, the SOM has a complexity of $O(N)$. In this work, we combine these two algorithms: first, we group the systematics vectors into SOM cells and then we further group the SOM cells into hierarchical clusters [^7] Training a small SOM in the first place instead of grouping big SOM into clusters will restrict the accuracy of the manifold’s mapping. Moreover, Moreover, It’s more flexible and faster to change the number of HC clusters from a pre-trained SOM than to change the dimension of a SOM and retrain.. We call this the “SOM+HC” method. After the training, the hierarchical clusters are projected back onto the sky, resulting in discrete regions occupied by galaxies with uniform systematics and selection effects. We then randomly redistribute galaxies across these regions and combine galaxies from all the clusters to form the OR. The details of the method are presented below.
@@
In this work, instead of training a SOM in the colour space, we train it in the “systematics space”, i.e., the data vectors are $N_{\mathrm{gal}}$ (number of galaxies) $N_{\mathrm{sys}}$-dimensional vectors, where $N_{\mathrm{sys}}$ is the number of systematics to account for ($N_{\mathrm{sys}}=5$ in our fiducial case). A SOM consists of $N_{\mathrm{dim}}\times N_{\mathrm{dim}}=N_{\mathrm{cell}}/,$ “cells” to represent the $N_{\mathrm{gal}}$ data vector. The positions of SOM cells in the $N_{\mathrm{sys}}$-dimensional systematics space are called “weight vectors”.
For the $n$-th galaxy, we use $\boldsymbol{V}_{n}$ to denote its systematics vector and $V_{n,i},i=1,2,...,N_{\mathrm{sys}}$ to denote its $i$-th component. We can calculate the Euclidean distance between the weight vector of each SOM cell $\boldsymbol{W}^{a},a=1,2,...,N{\mathrm{cell}}$ [^8] Note that a SOM is 2-dimensional, so one could use two cell indices (column and row indices). and the systematics vector as:
@@
    Initialise the SOM by setting up the initial position of $N_{\mathrm{cell}}$ weight vectors as the two highest principal component analysis (PCA) components. This is equivalent to initializing the weight vectors as the $N_{\mathrm{cell}}$ on the 2-dimensional subspace spanned by the first two eigenvectors of the correlation matrix of the data vectors. [^9] One can also initialise the SOM with random weight vectors.
@@
      Choose a data vector and update the weight vector as:
@@
    Iteratively merge two clusters separated by the shortest distance specified by the complete linkage criteria101010 In this work, we use the “complete linkage criteria” with Euclidean distance, for which the distance between clusters equals the Euclidean distance between two data vectors (one in each cluster) that are farthest away from each other.;
@@
    Cut the dendrogram where there are $NC$ clusters. See Figure 5 for an example of a dendrogram.
Figure 5: An example of a dendrogram which shows the clustering of 900 SOM cells into 20 HCs. The cells are clustered from the bottom to the top according to their Euclidean distance in the systematics space. The black dashed line shows the distance threshold where the cells are grouped into 20 clusters. SOM cells resulting in the same clusters are colour-coded with the same colour.
By construction, a galaxy cluster in the systematics space represents a particular combination of systematics. If the galaxy distribution in systematics space does not depend on the LSS of the Universe (we will validate this assumption in Section 5.2.2), we can assume that the number of galaxies in each cluster is determined by a uniform selection probability given by the synthetic systematics represented by that cluster. After being grouped into clusters through SOM+HC, the galaxies from one cluster are distributed into discrete regions in the survey footprint and are assumed to be uniformly depleted by the associated combination of systematics. The detailed process of generating OR from clustered galaxies is presented in the next subsection.
The number of hierarchical clusters $NC$ is an important parameter. If $NC$ is too large, there will be higher sampling noise in each cluster. In addition, high $NC$ will cause overfitting of the OR, resulting in over-correction in the 2PCF. If $NC$ is too small, the resolution in the systematics space will be too low to detect systematics variability. Therefore, we will evaluate the $NC$ value that optimises the trade-off between resolution and sampling noise in Section 5.

### v2

@@ 2.2 The measurement of two-point correlation function
In practice, the galaxy 2PCF is measured in angular bins by counting the number of galaxy pairs (one galaxy from catalogue $a$ and the other from $b$) with separation angles within each angular bin. Given a pair of galaxy catalogues (denoted as $D$ hereafter) and corresponding random catalogues (denoted as $R$ hereafter), the 2PCF can be measured using a standard Landy & Szalay (1993) estimator,
@@
where $\theta_{i}$ is the mean angle in the $i$-th angular bin. Here $(D^{a}D^{b})_{i}$ is the normalised number of galaxy pairs with separation angles within the $i$-th bin, which can be formally expressed as
@@
where $N_{D}$ is the number of galaxies, $\theta_{pq}$ is the separation angle between the $p$-th and $q-$th galaxies, $\Theta_{i}(\theta)$ is the rect function that equals 1 when $\theta$ falls in the $i-$th bin and 0 otherwise. $\delta_{\mathrm{K}}^{ab}$ is the Kronecker $\delta$ symbol. The denominator in Eq. (8) is the total number of galaxy pairs and the $\delta^{ab}$ term concerns removing galaxies paired with themselves for auto-correlation (i.e. $a=b$). The $RR$ term is likewise defined while $DR$ is defined without $\delta_{\mathrm{K}}^{ab}$.
For this work, instead of counting galaxy pairs in angular bins, we pixelised the galaxy distribution into the Healpix scheme (Gorski et al., 2005) and counted pixel pairs weighted by galaxy numbers in each pixel. This loses sub-pixel information, but we had three reasons for doing this: 1) pixelisation can speed up the calculation: the original KiDS catalogue contains $\sim 200$ million galaxies, while a pixelised galaxy map with Nside=2048 which we use throughout this work only has about one million non-zero pixels, so counting pixel pairs instead of galaxy pairs will significantly speed up the measurement; 2) as we introduce later, our method to recover ORs is pixel-based, so any sub-pixel variable depth cannot be corrected; 3) the pixel size for an Nside=2048 HEALPix map is 1.7 arcmin, corresponding to a physical size of 0.86 Mpc at $z=1.5$, which is much below the linear scales that we probe in this work.
@@
The random catalogue or map is used to factor out galaxy distributions induced by cosmology-unrelated systematics, including galactic and atmospheric foregrounds, instrument response, and survey strategy. If all those systematics are uniform across the survey footprint, the random catalogue is usually taken as a uniform Poisson random sample within the survey footprint with several times the mean galaxy number density. However, deep surveys, such as the KiDS survey, have non-uniform systematics that cause variable depth in the galaxy sample. A UR will fail to capture the spatial variation in galaxy number density caused by those selection effects and biases the 2PCF.
To de-bias the 2PCF, we can create a random catalogue with the same variable depth, namely an ‘organised random’ catalogue, or, for pixelised measurement, an OR weight map for the $R$ terms in the estimator. The method to recover the OR will be presented in Sect.. 4. We note that the OR can be used for summary statistics other than the angular 2PCF. We present the OR application in angular power spectra $C_{\ell}$ in Appendix B.
@@ 3.1 The Kilo-degree Survey and the Legacy data
Figure 1: Galaxy distribution of the KiDS-Legacy catalogue. The map is pixelised into HEALPix grids with $\texttt{Nside}=2048$. The map is colour-coded by galaxy number per pixel, which has a size of 1.7 arcmin. A tile-based selection pattern can be seen by eye. The shaded region in the colour bar shows the normalised distribution of galaxy number per pixel.
Table 1: Summary of the systematics of each galaxy that we used to recover the OR.
@@
 [^5] The second column indicates the spatial variance type: inter-tile (type A), intra-tile (type B), and tile-independent (type C). The other columns indicate their definitions, physical origins, and how they potentially select the galaxy sample.
@@
In this work, we use the KiDS-Legacy galaxy catalogue selected from the fifth data release (DR5) of the Kilo-degree Survey (KiDS; Wright et al., 2024). KiDS is a wide-field imaging survey that measures the positions and shapes of galaxies using the VLT Survey Telescope (VST) at the European Southern Observatory (ESO). Both the telescope and the survey were primarily designed for weak gravitational lensing applications. High-quality optical images are produced with VST-OmegaCAM, and these data are then combined with imaging from the VISTA Kilo-degree INfrared Galaxy survey (VIKING; Edge et al. 2013), allowing all sources in KiDS to be photometrically measured in nine optical and near-infrared bands: $ugriZYJHK_{\mathrm{s}}$ (Wright et al., 2019). Although the sky coverage of KiDS is smaller than some galaxy lensing surveys (such as DES, Abbott et al., 2016), galaxy photometric redshift estimation and redshift distribution calibration (especially at high redshift) benefit from complementary NIR information from VIKING (which was co-designed with KiDS to reach complementary depths in the NIR bands). Cosmological constraints have already been made available from DR3 and DR4 (Hildebrandt et al., 2017; Heymans et al., 2021). KiDS DR5 covers a survey area of 1347 $\mathrm{deg}^{2}$. It also includes an $i$-band re-observation of the full footprint, thereby increasing the effective $i$-band depth by 0.4 magnitudes and enabling multi-epoch science. There is a 27 $\mathrm{deg}^{2}$ overlap with deep spectroscopic surveys, which enables the robust calibration of photometric redshifts across the full survey footprint (Hildebrandt et al., 2021).
@@
The “KiDS-Legacy” sample used in this work is a subset of the full DR5 sample primarily determined by the availability of reliable shape measurements. It is the lensing sample that will be used for the fiducial KiDS DR5 cosmological analyses. Each galaxy in the KiDS-Legacy sample has ellipticities measured with the lensfit algorithm (Miller et al., 2013; Fenech Conti et al., 2017; Li et al., 2023a) for weak lensing analyses. The footprint of the KiDS DR5 is divided into a Northern and a Southern patch. Artefacts around bright, saturated starlight, planets, the Moon, satellite flares, aeroplanes, and higher-order reflections from very bright stars have been masked out, leaving an unmasked $\sim 1000\mathrm{deg}^{2}$ footprint. In addition, blended sources, unresolved binaries, transients, point sources flagged by lensfit, and sources with failed photometry, badly estimated shapes, low resolution, or zero lensing weight are removed from the full sample, resulting in a Legacy sample of 43,205,156 galaxies, corresponding to 12 galaxies per arcmin-2. A Bayesian photo-$z$ (BPZ, Benítez, 2000) estimation gives an approximated redshift range of $2.00.1<z_{\mathrm{B}}<$ 2.0. The redshift distribution is calibrated with a combination of SOM and clustering-redshift method(van den Busch et al., 2020), which will be presented in a companion paper. For more details about the construction of the KiDS-Legacy catalogue, we refer to Sect. 7 of Wright et al. (2024). The galaxy distribution is shown in Fig. 1. We can tell by eye that the galaxy distribution has a clear tile-like pattern that is not likely the LSS.
@@ 3. Type C,
A synthesis of systematics results in a complex selection pattern, namely variable depth, in the galaxy distribution. Figure 1 illustrates the galaxy map of the KiDS-Legacy catalogue, which exhibits a tile-based pattern of variable depth. For example, galaxies in the anomalous dark blue tile around ($\alpha=196\deg$, $\delta=1.5\deg$) in the centre of KiDS-North in Fig. 1 are significantly depleted. Heydenreich et al. (2020) studied the effect of variable depth on cosmic shear and concluded that the impact is insignificant for KiDS-like surveys. However, the variable depth will introduce significant bias in angular clustering measurement if not corrected. As galaxy clustering measurements form part of the KiDS-Legacy $3\times 2$pt and $6\times 2$pt analyses, it is essential to correct this bias.
@@
To correct the selection effects, we create an OR which reflects the same variable depth as the galaxy sample from the systematics given by the KiDS-Legacy catalogue. Those systematics generally originate from the Galactic foreground, atmospheric seeing, survey strategy, and instrument set-up, and are ideally independent of cosmology, galaxy properties, and the LSS. In the catalogue, systematics are described by metadata that are calculated or derived from the instrument set-up, observing conditions, image properties, and so on. In the following text, we use ‘systematics’ to also refer to the metadata we use to obtain the OR. We select five systematics that are most likely to have selection effects: {Level, PSF_size, PSF_ell, EXTINCTION_r, GAIA_nstar}. Their types of distribution, definitions, origins and selection effects are summarised in Table 5. These systematics, with the exception of GAIA_nstar, are measured in the $r$-band because this is the detection band and we do not consider tomographic redshift binning in this analysis.
@@
Figure 2 shows the maps of these systematics. Each systematics map is divided into south and north fields, colour-coded by the mean systematics value in each pixel. The shaded regions in the colour bars are the probability distributions of the systematics. From Fig. 2, we can see how Level, PSF_size, PSF_ell are distributed based on the $1^{\circ}\times 1^{\circ}$ tiles. Among them, Level and PSF_size are determined by the seeing of each exposure and are therefore relatively uniform in each tile, but vary from tile to tile; while PSF_ell depends on the curvature of the focal plane, which changes from the centre to the edge of each tile. On the other hand, EXTINCTION_r [^6] Given that the extinction in different bands is commonly just assumed to be the same template scaled by a different factor (thus they are fully correlated), we only use the $r$-band extinction for our model training. and Gaia_nstar reflect the spatial distribution of the Milky Way and are therefore more diffuse and independent of the tiles. To illustrate the selection effect of each systematics, we plot galaxy number contrast with respect to systematics values as black curves in the colour bars of Fig. 2. We note that the black curves share the same dynamic range between -0.25 and 0.25, so one can see that PSF_size has the strongest selection effect. In addition, the selection effects of EXTINCTION_r and Gaia_nstar are slightly non-linear.
@@
It should be noted that we could recover the OR with all the systematics such as PSF shape components and background counts, but in practice, it is advantageous to only use a subset of systematics that is most likely to select galaxies to reduce computation time. The rest of the systematics are either not correlated with galaxy numbers or are strongly correlated with the selected systematics. In Appendix C we run our method with all the systematics provided in the catalogue and prove that the performance does not improve. In the following sections of this paper, we present and test the method for recovering the organised random from these five systematics.
@@
Figure 2: Maps of the KiDS-Legacy systematic considered in this paper. Each map is divided into northern and southern fields, plotted together with their colour bars. The black curves over-plotted in the colour bars show the relationship between galaxy contrast and systematics value, and the dynamic ranges are $[-0.25,0.25]$. The shaded regions show the probability distributions of each systematics.
@@ 4 The SOM+HC method to recover organised randoms
Figure 3: Flowchart of the measurement of the 2PCF with reconstructed organised randoms. The colour block after the 2PCF indicates the colours shown in the following $w(\theta)-\theta$ figures.
Figure 3 shows a flowchart of the 2PCF measurement of galaxies with variable depth. Through observation, we obtain a depleted catalogue of galaxies with systematics evaluated for each galaxy. If we follow the upper half of the flowchart by using UR to compute the 2PCF, we get a biased “UR” 2PCF. Instead, we can recover OR to correct for non-uniform selection effects and obtain the unbiased “OR” 2PCF. In this section we introduce the method that we used to generate OR, namely a combination of self-organising map (SOM) and hierarchical clustering (HC). Figure 4 summarises this “SOM+HC” method.
Figure 4: Flowchart illustrating the SOM+HC method to recover OR and correct selection effects in the 2PCF. Starting from the top: 1. Input systematics; 2. SOM training: left panel: SOM grid colour-coded by galaxy number in each cell; right panel: projection of systematics vectors (small grey dots) and SOM cells (red dots connecting) on the plane of two systematics. The projected adjacent SOM cells are connected with black lines; 3. HC output: left panel: SOM cells colour-coded according to hierarchical cluster indices; right panel: systematics vectors and weight vectors colour-coded by corresponding cluster indices; 4. Effective areas corresponding to galaxies from each cluster; 5. Recovered OR weight map which will be used in a subsequent 2PCF measurement.
@@ 4.1 The selection effects induced by systematics
If the systematics are uncorrelated, the overall selection function is a multiplication of the selection functions of individual systematics. One can reconstruct the total selection function by modelling the selection effect of each systematics (Rodríguez-Monroy et al., 2022). In practice, however, it is not always possible to construct reliable quantitative models to describe the effect of each systematics on the galaxy number density. In addition, systematics can be correlated (e.g. extinction in different bands, extinction and GAIA stars), making the overall selection function more complicated.
The key to recovering the ORs is to group galaxies into subsamples, each sharing a similar selection effect. Galaxies from each subsample occupy a subregion in the survey footprint that is presumably selected uniformly. Therefore, we can generate a UR in the same sky region for each group, and combine them to obtain the OR. In this work, we use a combination of hierarchical clustering (Murtagh & Contreras, 2012, HC) and SOM, namely the SOM+HC method, to group the galaxies and recover the OR weight to account for the selection effect. This method does not assume formal models for selection functions and their correlations, so it is flexible enough to account for arbitrarily complex selection effects.
@@ 4.2 Self-organising map and hierarchical clustering
HC is a widely used clustering algorithm. It has the advantages of flexibility, robustness, and interpretability, but is computationally expensive (the computational time scales with the data volume as $O(N^{2})$, where $N$ is the number of galaxies). On the other hand, the SOM has a complexity of $O(N)$. In this work, we combine these two algorithms: first, we group the systematics vectors into SOM cells and then we further group the SOM cells into hierarchical clusters [^8] Training a small SOM in the first place instead of grouping big SOM into clusters will restrict the accuracy of the manifold’s mapping. Moreover, It’s more flexible and faster to change the number of HC clusters from a pre-trained SOM than to change the dimension of a SOM and retrain.. After the training, the hierarchical clusters are projected back onto the sky, resulting in discrete regions occupied by galaxies with uniform systematics and selection effects. We then randomly redistribute galaxies across these regions and combine galaxies from all the clusters to form the OR. The details of the method are presented below.
@@
In this work, instead of training a SOM in the colour space, we train it in the systematics space: the data vectors are $N_{\mathrm{gal}}$ (number of galaxies) $N_{\mathrm{sys}}$-dimensional vectors, where $N_{\mathrm{sys}}$ is the number of systematics to account for ($N_{\mathrm{sys}}=5$ in our fiducial case). A SOM consists of $N_{\mathrm{dim}}\times N_{\mathrm{dim}}=N_{\mathrm{cell}}/,$ cells to represent the $N_{\mathrm{gal}}$ data vector. The positions of SOM cells in the $N_{\mathrm{sys}}$-dimensional systematics space are called weight vectors.
For the $n$-th galaxy, we use $\boldsymbol{V}_{n}$ to denote its systematics vector and $V_{n,i},i=1,2,...,N_{\mathrm{sys}}$ to denote its $i$-th component. We can calculate the Euclidean distance between the weight vector of each SOM cell $\boldsymbol{W}^{a},a=1,2,...,N_{\mathrm{cell}}$ [^9] A SOM is two-dimensional, so one could use two cell indices (column and row indices). and the systematics vector as
@@
    Initialise the SOM by setting up the initial position of $N_{\mathrm{cell}}$ weight vectors as the two highest principal component analysis (PCA) components. This is equivalent to initialising the weight vectors as the $N_{\mathrm{cell}}$ on the 2-dimensional subspace spanned by the first two eigenvectors of the correlation matrix of the data vectors. [^10] One can also initialise the SOM with random weight vectors.
@@
      Choose a data vector and update the weight vector as
@@
    Iteratively merge two clusters separated by the shortest distance specified by the complete linkage criteria111111 In this work, we use the “complete linkage criteria” with Euclidean distance, for which the distance between clusters equals the Euclidean distance between two data vectors (one in each cluster) that are farthest away from each other.;
@@
    Cut the dendrogram where the cells are merged into the desired number of clusters (we use $NC$ to denote the number of clusters hereafter). See Fig. 5 for an example of a dendrogram.
Figure 5: Example of a dendrogram showing the clustering of 900 SOM cells into 20 HCs. The cells are clustered from the bottom to the top according to their Euclidean distance in the systematics space. The black dashed line shows the distance threshold where the cells are grouped into 20 clusters. SOM cells resulting in the same clusters are colour-coded with the same colour.
By construction, a galaxy cluster in the systematics space represents a particular combination of systematics. If the galaxy distribution in systematics space does not depend on the LSS of the Universe (we validate this assumption in Sect. 5.2.2), we can assume that the number of galaxies in each cluster is determined by a uniform selection probability given by the synthetic systematics represented by that cluster. After being grouped into clusters through SOM+HC, the galaxies from one cluster are distributed into discrete regions in the survey footprint and are assumed to be uniformly depleted by the associated combination of systematics. The detailed process of generating OR from clustered galaxies is presented in the next subsection.
The number of hierarchical clusters $NC$ is an important parameter. If $NC$ is too large, there will be higher sampling noise in each cluster. In addition, high $NC$ will cause over-fitting of the OR, resulting in over-correction in the 2PCF. If $NC$ is too small, the resolution in the systematics space will be too low to detect systematics variability. Therefore, we evaluate the $NC$ value that optimises the trade-off between resolution and sampling noise in Sect. 5.